4x+2y=-10 -2x+3y=33
Answer by stanbon(75887)

4x+2y=-10
-2x+3y=33
--------------
2x + y = -5
-2x + 3y = 33
----------------------
Add and solve for "y":
4y = 28
y = 7
-----
Solve for "x":
2x + y = -5
2x = -12
x = -6
------------
Cheers,
Stan H.
-----------